The braking distance d of a car traveling v mi/hr is approximated by d=v+(v^2/20). Determine velocities that result in braking distance of less than 5ft.
1
Expert's answer
2020-12-02T12:26:02-0500

To begin with, we need to recognize given valuables:

d=5 ft = 0,00094697 mi and also the equation which connects velocity and distance.

d=v+(v2/20).

So if we put the given length or given breaking distance into equation, we will find the velocity that the car can stop in less than 5 ft.

0,00094697 mi = v+("\frac{v^2}{20}"),

v2+20"\times" v - 20"\times" 0.00094697=0

we will solve this equation by finding its roots:

but first we need to compute discriminant then we can find roots:

D=b2-4"\times a \times c" ,

D=202+4"\times 1\times0.0189394" =400.07557

there are two roots and one of them is negative and another of them is positive. We will find positive one:

v="\frac{-20+\sqrt{\smash[b]{D}}}{2}" =0.000946 "\frac{mi}{hr}"

if we compute previous equation, we will reach:

v=0.000946 "\frac{mi}{hr}" .

if the car move with v then it can stop less than 5 ft.